# Supersymmetry and Localization on Three-Dimensional Orbifolds

### v1

@@ 5.1.1 General Construction
$=\frac{\prod_{j_{2}}{\left[{-p{\left({n_{N}j_{2}-t_{N}\mathfrak{m}_{2}}\right)}-\mathfrak{s}_{w}q{\left({-n_{S}j_{2}+t_{S}\mathfrak{m}_{2}}\right)}-{\rm i}\mathfrak{s}_{w}pq\sigma_{c}+\frac{q_{R}^{B}}{2}{\left({p\alpha_{2}+q\mathfrak{s}_{w}\alpha_{3}}\right)}}\right]}}{\prod_{m_{1},n_{1}}{\left[{p{\left({n_{N}j_{1}-t_{N}\mathfrak{m}_{1}}\right)}+\mathfrak{s}_{w}q{\left({-n_{S}j_{1}+t_{S}\mathfrak{m}_{1}}\right)}+{\rm i}\mathfrak{s}_{w}pq\sigma_{c}-\frac{q_{R}^{\phi}}{2}{\left({p\alpha_{2}+q\mathfrak{s}_{w}\alpha_{3}}\right)}}\right]}}~{},$
@@
$=\frac{\prod_{j_{2}}-\mathfrak{s}_{w}{\left({\mathfrak{s}_{w}pn_{N}-qn_{S}}\right)}j_{2}+\mathfrak{s}_{w}{\left({\mathfrak{s}_{w}pt_{N}-qt_{S}}\right)}\mathfrak{m}_{2}-{\rm i}\mathfrak{s}_{w}pq\sigma_{c}+q_{R}^{B}{\left[{{\left({pn_{N}-q\mathfrak{s}_{w}n_{S}}\right)}c_{R}+{\left({q\mathfrak{s}_{w}t_{S}-pt_{N}}\right)}\lambda_{R}}\right]}}{\prod_{j_{1}}\mathfrak{s}_{w}{\left({\mathfrak{s}_{w}pn_{N}-qn_{S}}\right)}j_{1}-\mathfrak{s}_{w}{\left({\mathfrak{s}_{w}pt_{N}-qt_{S}}\right)}\mathfrak{m}_{1}+{\rm i}\mathfrak{s}_{w}pq\sigma_{c}-q_{R}^{\phi}{\left[{{\left({pn_{N}-q\mathfrak{s}_{w}n_{S}}\right)}c_{R}+{\left({q\mathfrak{s}_{w}t_{S}-pt_{N}}\right)}\lambda_{R}}\right]}}~{},$
$=\frac{\prod_{j_{2}}{\left[{-\alpha_{\rm 2d}{\left({j_{2}-q_{R}^{B}c_{R}}\right)}-{\left({\mathfrak{m}_{2}-q_{R}^{B}\lambda_{R}}\right)}-\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}}\right]}}{\prod_{j_{1}}{\left[{\alpha_{\rm 2d}{\left({j_{1}-q_{R}^{\phi}c_{R}}\right)}+{\left({\mathfrak{m}_{1}-q_{R}^{\phi}\lambda_{R}}\right)}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}}\right]}}~{},$ (5.27)
where we used
$\mathfrak{s}_{w}n_{N}p-n_{S}q=\alpha_{\rm 2d}{\left({t_{S}q-\mathfrak{s}_{w}t_{N}p}\right)}~{}.$ (5.28)
Altogether,
$Z_{\text{1-L}}$ $=\frac{\prod_{j_{2}}{\left[{-\alpha_{\rm 2d}{\left({j_{2}-q_{R}^{B}c_{R}}\right)}-{\left({\mathfrak{m}_{2}-q_{R}^{B}\lambda_{R}}\right)}-\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}}\right]}}{\prod_{j_{1}}{\left[{\alpha_{\rm 2d}{\left({j_{1}-q_{R}^{\phi}c_{R}}\right)}+{\left({\mathfrak{m}_{1}-q_{R}^{\phi}\lambda_{R}}\right)}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}}\right]}}~{},$
$m_{i}$ $=n_{N}j_{i}-t_{N}\mathfrak{m}_{i}~{},\qquad n_{i}=-n_{S}j_{i}+t_{S}\mathfrak{m}_{i}~{},$
$\lambda_{R}$ $=-\frac{\mathfrak{s}_{w}n_{N}+n_{S}}{2}~{},\qquad c_{R}=-\frac{t_{S}+\mathfrak{s}_{w}t_{N}}{2}=\frac{\mathfrak{s}_{w}n_{N}-n_{S}}{4\,n_{N}n_{S}}+\frac{\lambda_{R}{\left({n_{N}t_{S}+n_{S}t_{N}}\right)}}{2\,n_{N}n_{S}}~{},$
$m_{1}$ $=\ell_{1}~{},\qquad n_{1}=\mathfrak{s}_{w}k_{1}~{},\qquad m_{2}=-\ell_{2}~{},\qquad n_{2}=-\mathfrak{s}_{w}k_{2}~{},\qquad\ell_{1,2},k_{1,2}\in\mathbb{N}~{}.$ (5.29)
5.1.2 Twisted Ellipsoid
The ellipsoid with a non-trivial R-symmetry flux is obtained by setting
$\mathfrak{s}_{w}=1~{},$ $n_{N}=n_{S}=1~{},\qquad t_{S}=1+t_{N}~{},$ (5.30)
implying
$m_{i}$ $=j_{i}-t_{N}\mathfrak{m}_{i}~{},\qquad n_{i}=-j_{i}+t_{N}\mathfrak{m}_{i}+\mathfrak{m}_{i}~{},$
$\lambda_{R}$ $=-1~{},\qquad c_{R}=-\frac{2t_{N}+1}{2}~{},$
$m_{1}$ $=\ell_{1}~{},\qquad n_{1}=k_{1}~{},\qquad m_{2}=-\ell_{2}~{},\qquad n_{2}=-k_{2}~{},\qquad\ell_{1,2},k_{1,2}\in\mathbb{N}~{},$ (5.31)
which give the intervals
$t_{N}\mathfrak{m}_{1}\leq j_{1}\leq t_{S}\mathfrak{m}_{1}=t_{N}\mathfrak{m}_{1}+\mathfrak{m}_{1}~{},$ $t_{N}\mathfrak{m}_{2}+\mathfrak{m}_{2}=t_{S}\mathfrak{m}_{2}\leq j_{2}\leq t_{N}\mathfrak{m}_{2}~{}.$ (5.32)
The one-loop determinant can be put in the form reported in [36] via the shifts
$j_{1}=i_{1}+\frac{\mathfrak{m}_{1}}{2}{\left({t_{N}+t_{S}}\right)}=i_{1}+\mathfrak{m}_{1}{\left({t_{N}+\frac{1}{2}}\right)}~{},$
$j_{2}=i_{2}+\frac{\mathfrak{m}_{2}}{2}{\left({t_{N}+t_{S}}\right)}=i_{2}+\mathfrak{m}_{2}{\left({t_{N}+\frac{1}{2}}\right)}~{},$ (5.33)
with $i_{1,2}$ being half-integers satisfying
$-\frac{\mathfrak{m}_{1}}{2}\leq i_{1}\leq\frac{\mathfrak{m}_{1}}{2}~{},$ $\frac{\mathfrak{m}_{2}}{2}\leq i_{2}\leq-\frac{\mathfrak{m}_{2}}{2}~{}.$ (5.34)
The integers $\mathfrak{m}_{1,2}$ are the fluxes assigned to $\phi,B$ on the ellipsoid, respectively [33] . Hence,
$\mathfrak{m}_{1}$ $=q_{G}\mathfrak{m}+q_{R}^{\phi}\mathfrak{f}_{R}=\mathfrak{m}+q_{R}^{\phi}\lambda_{R}=\mathfrak{m}-r=-\mathfrak{r}~{},$
$\mathfrak{m}_{2}$ $=q_{G}\mathfrak{m}+q_{R}^{B}\mathfrak{f}_{R}=\mathfrak{m}+q_{R}^{B}\lambda_{R}=\mathfrak{m}+2-r=2-\mathfrak{r}~{},$ (5.35)
giving
$\frac{\mathfrak{r}}{2}\leq i_{1}\leq-\frac{\mathfrak{r}}{2}~{},$ $-\frac{\mathfrak{r}-2}{2}\leq i_{2}\leq\frac{\mathfrak{r}-2}{2}~{}.$ (5.36)
Finally,
$Z_{\text{1-L}}|_{\mathfrak{r}\geq 2}$ $=\prod_{i_{2}={\left({2-\mathfrak{r}}\right)}/2}^{{\left({\mathfrak{r}-2}\right)}/2}{\left({-\alpha_{\rm 2d}i_{2}-\gamma_{G}}\right)}={\left({-}\right)}^{\mathfrak{r}-1}\prod_{i_{2}={\left({2-\mathfrak{r}}\right)}/2}^{{\left({\mathfrak{r}-2}\right)}/2}{\left({\alpha_{\rm 2d}i_{2}+\gamma_{G}}\right)}~{},$
$Z_{\text{1-L}}|_{\mathfrak{r}=1}$ $=1~{},$
$Z_{\text{1-L}}|_{\mathfrak{r}\leq 0}$ $=\prod_{i_{1}=\mathfrak{r}/2}^{-\mathfrak{r}/2}{\left({\alpha_{\rm 2d}i_{2}+\gamma_{G}}\right)}^{-1}~{},$ (5.37)
matching [36] . The gauge fugacity reads
$\gamma_{G}=\alpha_{\rm 2d}\frac{\mathfrak{m}}{2}{\left({t_{N}+t_{S}}\right)}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}+\mathfrak{m}~{}.$ (5.38)
The argument of the product in $Z_{\text{1-L}}$ is independent of the R-charge; namely, the R-symmetry fugacity $\gamma_{R}$ is vanishing. This is compatible with the 3d constraint ${\left({2\gamma_{R}}\right)}\in\mathbb{Z}$ valid for a $\mathcal{N}=2$ theory on $S^{2}\times S^{1}$ with R-symmetry flux through $S^{2}$.
5.1.3 Untwisted Ellipsoid
The ellipsoid with a no R-symmetry flux is obtained by setting
$\mathfrak{s}_{w}=-1~{},$ $n_{N}=n_{S}=1~{},\qquad t_{S}=1+t_{N}~{},$ (5.39)
leading to
$m_{i}$ $=n_{N}j_{i}-t_{N}\mathfrak{m}_{i}~{},\qquad n_{i}=-n_{S}j_{i}+t_{S}\mathfrak{m}_{i}~{},$
$\lambda_{R}$ $=0~{},\qquad c_{R}=-\frac{1}{2}~{},$
$m_{1}$ $=\ell_{1}~{},\qquad n_{1}=-k_{1}~{},\qquad m_{2}=-\ell_{2}~{},\qquad n_{2}=k_{2}~{},\qquad\ell_{1,2},k_{1,2}\in\mathbb{N}~{}.$ (5.40)
The integers $\mathfrak{m}_{1,2}$ are the fluxes assigned to $\phi,B$ on the ellipsoid, respectively [33] . Hence,
$\mathfrak{m}_{1}$ $=q_{G}\mathfrak{m}=\mathfrak{m}=\mathfrak{m}_{2}~{}.$ (5.41)
We then obtain the intervals
$j_{1}\geq t_{N}\mathfrak{m}~{},$ $j_{1}\geq t_{S}\mathfrak{m}~{},$
$j_{2}\leq t_{N}\mathfrak{m}~{},$ $j_{2}\leq t_{S}\mathfrak{m}~{}.$ (5.42)
The ranges above would introduce $|\mathfrak{m}|$ in the one-loop determinant. We can avoid that by picking up ranges from dual poles, which is the prescription employed in the index theorem computation. Therefore, if
$j_{1}\geq t_{N}\mathfrak{m}~{},$ $j_{2}\leq t_{S}\mathfrak{m}~{},$ (5.43)
@@ 5.1.3 Untwisted Ellipsoid
$j_{1}=t_{N}\mathfrak{m}+i=\frac{\mathfrak{m}}{2}{\left({t_{N}+t_{S}}\right)}-\frac{\mathfrak{m}}{2}+i~{},$
$j_{2}=t_{S}\mathfrak{m}-i=\frac{\mathfrak{m}}{2}{\left({t_{N}+t_{S}}\right)}+\frac{\mathfrak{m}}{2}-i~{},\qquad i\in\mathbb{N}~{},$ (5.44)
we find
$Z_{\text{1-L}}=\prod_{i\in\mathbb{N}}\frac{i+1-\frac{r}{2}-\frac{\mathfrak{m}}{2}+{\rm i}\mathfrak{u}}{i+\frac{r}{2}-\frac{\mathfrak{m}}{2}-{\rm i}\mathfrak{u}}~{},$ (5.45)
matching [33]. The gauge fugacity is
$\gamma_{G}=\alpha_{\rm 2d}\frac{\mathfrak{m}}{2}{\left({t_{N}+t_{S}}\right)}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}+\mathfrak{m}=-{\rm i}\alpha_{\rm 2d}\mathfrak{u}~{},$ (5.46)
and R-symmetry fugacity
$\gamma_{R}=\frac{1}{2}\alpha_{\rm 2d}=\alpha_{\rm 2d}\frac{\chi{\left({S^{2}}\right)}}{4}~{},$ (5.47)
which is compatible with the 3d constraint valid for an anti-twisted $\mathcal{N}=2$ theory on $\mathbb{\Sigma}\times S^{1}$.
5.1.4 Twisted Spindle
The twisted spindle is obtained by setting
$\mathfrak{s}_{w}=1~{},$ (5.48)
with $n_{N}$ and $n_{S}$ being positive integers fulfilling $\gcd{\left({n_{N},n_{S}}\right)}=1$. Therefore,
$m_{i}$ $=n_{N}j_{i}-t_{N}\mathfrak{m}_{i}~{},\qquad n_{i}=-n_{S}j_{i}+t_{S}\mathfrak{m}_{i}~{},$
$\lambda_{R}$ $=-\frac{n_{N}+n_{S}}{2}~{},\qquad c_{R}=-\frac{t_{S}+t_{N}}{2}=\frac{1}{4}{\left({\frac{1}{n_{S}}-\frac{1}{n_{N}}}\right)}+\frac{\lambda_{R}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}~{},$
$m_{1}$ $=\ell_{1}~{},\qquad n_{1}=k_{1}~{},\qquad m_{2}=-\ell_{2}~{},\qquad n_{2}=-k_{2}~{},\qquad\ell_{1,2},k_{1,2}\in\mathbb{N}~{}.$ (5.49)
These correspond to the intervals
${\left\lceil{\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}}\right\rceil}\leq j_{1}\leq{\left\lfloor{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}}\right\rfloor}~{},$ ${\left\lceil{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rceil}\leq j_{2}\leq{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}~{}.$ (5.50)
By generalizing [33], we have
$\mathfrak{m}_{1}$ $=q_{G}\mathfrak{m}+q_{R}^{\phi}\mathfrak{f}_{R}=\mathfrak{m}+q_{R}^{\phi}\lambda_{R}=\mathfrak{m}-\frac{r}{2}{\left({n_{N}+n_{S}}\right)}~{},$
$\mathfrak{m}_{2}$ $=q_{G}\mathfrak{m}+q_{R}^{B}\mathfrak{f}_{R}=\mathfrak{m}+q_{R}^{B}\lambda_{R}=\mathfrak{m}+\frac{2-r}{2}{\left({n_{N}+n_{S}}\right)}~{},$ (5.51)
giving the partition function
$Z_{\text{1-L}}|_{\mathfrak{m}_{2}\leq 0}=\prod_{j_{2}={\left\lceil{{t_{S}\mathfrak{m}_{2}}/{n_{S}}}\right\rceil}}^{{\left\lfloor{{t_{N}\mathfrak{m}_{2}}/{n_{N}}}\right\rfloor}}{\left[{-\alpha_{\rm 2d}{\left({j_{2}-q_{R}^{B}c_{R}}\right)}-\mathfrak{m}-\frac{{\rm i}pq\sigma_{c}}{t_{S}q-t_{N}p}}\right]}~{},$
$Z_{\text{1-L}}|_{1\leq{\left({r/2}\right)}{\left({n_{N}+n_{S}}\right)}-\mathfrak{m}\leq n_{N}+n_{S}-1}=1~{},$
$Z_{\text{1-L}}|_{\mathfrak{m}_{1}\geq 0}=\prod_{j_{1}={\left\lceil{{t_{N}\mathfrak{m}_{1}}/{n_{N}}}\right\rceil}}^{{\left\lfloor{{t_{S}\mathfrak{m}_{1}}/{n_{S}}}\right\rfloor}}{\left[{\alpha_{\rm 2d}{\left({j_{1}-q_{R}^{\phi}c_{R}}\right)}+\mathfrak{m}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-t_{N}p}}\right]}^{-1}~{}.$ (5.52)
The same result can be rewritten in another form by noticing that
$j_{i}$ $=t_{N}n_{i}+t_{S}m_{i}~{},$
$=\frac{b_{i}}{2n_{N}n_{S}}+\frac{\mathfrak{m}_{i}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}\in\mathbb{Z}~{},$ (5.53)
where $b_{i}\in\mathbb{Z}$ is such that
$b_{i}=n_{S}m_{i}-n_{N}n_{i}~{},$ $n_{S}m_{i}+n_{N}n_{i}=\mathfrak{m}_{i}~{},$
$-\mathfrak{m}_{1}\leq b_{1}\leq\mathfrak{m}_{1}~{},$ $\mathfrak{m}_{2}\leq b_{2}\leq-\mathfrak{m}_{2}~{},$ (5.54)
yielding
$Z_{\text{1-L}}|_{\mathfrak{m}_{2}\leq 0}=\prod_{b_{2}=\mathfrak{m}_{2}}^{-\mathfrak{m}_{2}}{\left\{{-\alpha_{\rm 2d}{\left[{\frac{b_{2}}{2n_{N}n_{S}}-q_{R}^{B}{\left({\frac{n_{N}-n_{S}}{4\,n_{N}n_{S}}}\right)}}\right]}-\gamma_{G}}\right\}}~{},$
$Z_{\text{1-L}}|_{1\leq{\left({r/2}\right)}{\left({n_{N}+n_{S}}\right)}-\mathfrak{m}\leq n_{N}+n_{S}-1}=1~{},$
$Z_{\text{1-L}}|_{\mathfrak{m}_{1}\geq 0}=\prod_{b_{1}=-\mathfrak{m}_{1}}^{\mathfrak{m}_{1}}{\left\{{\alpha_{\rm 2d}{\left[{\frac{b_{1}}{2n_{N}n_{S}}-q_{R}^{\phi}{\left({\frac{n_{N}-n_{S}}{4\,n_{N}n_{S}}}\right)}}\right]}+\gamma_{G}}\right\}}^{-1}~{},$ (5.55)
where the gauge fugacity is
$\gamma_{G}=\mathfrak{m}+\alpha_{\rm 2d}\frac{\mathfrak{m}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-t_{N}p}~{},$ (5.56)
while the R-symmetry fugacity reads
$\gamma_{R}=\alpha_{\rm 2d}{\left({\frac{n_{N}-n_{S}}{4\,n_{N}n_{S}}}\right)}=\frac{\alpha_{\rm 2d}}{4}\xi~{},$ $\xi=\frac{1}{n_{S}}-\frac{1}{n_{N}}~{},$ (5.57)
where the latter is consistent with the two-dimensional reduction of the three-dimensional constraint for a $\mathcal{N}=2$ theory on $\mathbb{\Sigma}\times S^{1}$ with R-symmetry twist on $\mathbb{\Sigma}$.
In fact, if we introduce $i_{2}$ by
$j_{2}=-i_{2}+{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}~{},$ $0\leq i_{2}\leq{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}-{\left\lceil{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rceil}~{},$ (5.58)

### v2

@@ 5.2 Topologically twisted spindle
$Z_{\rm 2d}^{\mathcal{B}}$ $=\prod_{\ell^{(1)}=1+{\left\lfloor{{\mathfrak{p}_{1}}/{n_{1}}}\right\rfloor}}^{-1-{\left\lfloor{-{\mathfrak{p}_{2}}/{n_{2}}}\right\rfloor}}{\left({{\omega}_{\rm 2d}\,\ell^{(1)}+\mathfrak{m}+{\rm i}\,\sigma_{\rm 2d}}\right)}~{}.$ (5.40)
Both $Z_{\rm 2d}^{\Phi}$ and $Z_{\rm 2d}^{\mathcal{B}}$ are encoded by the formula
@@
$Z_{\rm 2d}^{\rm CM}=\prod_{\ell\in\mathbb{N}}\frac{{\omega}_{\rm 2d}{\left({\ell+1+{\left\lfloor{-\mathfrak{p}_{2}/n_{2}}\right\rfloor}}\right)}-\mathfrak{m}-{\rm i}\sigma_{\rm 2d}}{{\omega}_{\rm 2d}{\left({\ell-{\left\lfloor{\mathfrak{p}_{1}/n_{1}}\right\rfloor}}\right)}-\mathfrak{m}-{\rm i}\sigma_{\rm 2d}}~{},$ (5.41)
which, after relabelling ${\left({\mathfrak{m}+{\rm i}\sigma_{\rm 2d}}\right)}\to{\left({-\varphi_{G}}\right)}$ and
${\left({t_{1},n_{1}}\right)}\to{\left({{t_{+}},{n_{+}}}\right)}~{},$ ${\left({t_{2},n_{2}}\right)}\to{\left({{t_{-}},{n_{-}}}\right)}~{},$
${\left({t_{1}\mathfrak{m},\mathfrak{p}_{1}}\right)}\to{\left({{\mathfrak{m}_{+}},\mathfrak{p}_{+}}\right)}~{},$ ${\left({t_{2}\mathfrak{m},\mathfrak{p}_{2}}\right)}\to{\left({{\mathfrak{m}_{-}},\mathfrak{p}_{-}}\right)}~{},$ (5.42)
precisely corresponds to (3.83) for ${\upsigma}=+1$, which represents the one-loop determinant of a chiral multiplet coupled to an Abelian vector multiplet on a topologically twisted spindle. Furthermore, setting ${n_{+}}={n_{-}}=+1$ and following the methodology detailed in Section 3 gives the chiral-multiplet one-loop determinant on an A-twisted $S^{2}$, as calculated in [37].
5.3 Anti-twisted spindle
If ${\mathfrak{s}_{\omega}}=-1$, then
$c_{R}=\frac{1}{2}{\left({t_{1}-t_{2}}\right)}~{},$ $\lambda_{R}=\frac{1}{2}{\left({n_{1}-n_{2}}\right)}~{},$ (5.43)
@@ 5.3 Anti-twisted spindle
$-n_{2}j^{(1)}+t_{2}\mathfrak{f}^{(1)}\geq 0~{},$ $n_{1}j^{(1)}-t_{1}\mathfrak{f}^{(1)}\leq 0~{},$
$-n_{2}j^{(2)}+t_{2}\mathfrak{f}^{(2)}\leq 0~{},$ $n_{1}j^{(2)}-t_{1}\mathfrak{f}^{(2)}\geq 0~{},$ (5.44)
meaning
$j^{(1)}\leq{\left\lfloor{t_{2}\mathfrak{f}^{(1)}/n_{2}}\right\rfloor}~{},$ $j^{(1)}\leq{\left\lfloor{t_{1}\mathfrak{f}^{(1)}/n_{1}}\right\rfloor}~{},$
$j^{(2)}\geq{\left\lceil{t_{2}\mathfrak{f}^{(2)}/n_{2}}\right\rceil}~{},$ $j^{(2)}\geq{\left\lceil{t_{1}\mathfrak{f}^{(2)}/n_{1}}\right\rceil}~{}.$ (5.45)
The inequalities above would lead to the inclusion of absolute values of fluxes in the one-loop determinant. This can be circumvented by selecting ranges from dual poles, as prescribed by the index theorem. Indeed, if
$j^{(1)}$ $\leq{\left\lfloor{t_{2}\mathfrak{f}^{(1)}/n_{2}}\right\rfloor}=-1-{\left\lfloor{-\mathfrak{p}_{2}/n_{2}}\right\rfloor}+{\left({r-2}\right)}c_{R}~{},$
$j^{(2)}$ $\geq{\left\lceil{t_{1}\mathfrak{f}^{(2)}/n_{1}}\right\rceil}={\left\lceil{\mathfrak{p}_{1}/n_{1}}\right\rceil}+r\,c_{R}~{},$ (5.46)